### v3

# AI-Powered CPS-Enabled Vulnerable-User-Aware Urban Transportation Digital Twin: Methods and Applications

Abstract
We present methods and applications for the development of digital twins (DT) for urban traffic management. While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived. In order to add value to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies, in other words, cyberphysical systems. This paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
footnotetext: This work has been published in IEEE Transactions on Intelligent Transportation Systems (IEEE T-ITS).footnotetext: Yongjie Fu, Mehmet K.Turkcan, Zhaobin Mo, and Xuan Di are with the Department of Civil Engineering and Engineering Mechanics, Columbia University. Emails: {yf2578, mkt2126, zm2302, sharon.di}@columbia.edufootnotetext: Mehmet K. Turkcan and Mahshid Ghasemi contributed equally to this work.footnotetext: Xuan Di, Zoran Kostic, and Gil Zussman are also with the Data Science Institute, Columbia University.footnotetext: Mahshid Ghasemi, Chengbo Zang, Abhishek Adhikari, Zoran Kostic, and Gil Zussman are with the Department of Electrical Engineering, Columbia University. Emails: {mg4089, cz2678, abhishek.adhikari, zk2172, gil.zussman}@columbia.edu
Keywords: Digital twin, AI, Urban traffic management
1 Introduction
Urban transportation systems are complex to model and simulate, due to heterogeneous road users (such as cars, pedestrians, cyclists, scooters) interacting in multimodal traffic environments consisting of public and private travel modes. With fast-changing traffic evolution in time and space, traffic simulation, if improperly calibrated, might produce traffic management strategies that largely deviate from the reality, potentially leading to suboptimal or even detrimental outcomes. With ubiquitous sensors in smart cities, it is the time to augment conventional traffic simulators, many of which were developed in an era when only “small data” became available. Emerging traffic sensors are expected to generate big volumes of data, transmitted via communication networks and processed on edge cloud computing with artificial intelligence (AI) for real-time traffic management. Such a transformation calls for the development of a new paradigm, namely, digital twin (DT), which will push the envelope in urban transportation management.
Literally, DT is the digital replica of a physical object or asset [73], where a digital world mirrors a physical world for real-time diagnosis, prognosis, and decision making. Recent years have seen a growing amount of studies on DT in various domains [73, 110, 118], including a sizable body of articles on vehicular DTs [179, 215, 84, 260, 114].
With recent explosive growth of literature on DTs, we would like to restrict the scope of this paper to applications in the urban setting, especially when vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [199]) are an integral part of the system and also potential users of the DT. The studies of DTs for urban traffic management, especially involving VRUs, are lacking, partly because the development of a DT for a system is non-trivial, particularly when involved with humans. This paper presents methods and applications for the development of DTs for urban transportation systems. We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods.
The rest of this paper is organized below. Section 2 introduces transportation DTs enabled by cyberphysical systems, and position this paper; Sections 3 and 4 review each component within CPS, and use cases for urban transportation, respectively. Section 5 demonstrates the available testbeds for DT testing and validation. Section 6 concludes our work, presents potential research directions and open questions.
2 Preliminaries
The definition of DT has evolved rapidly. Despite presenting its own version, articles share common elements and resemble certain characteristics [73, 152, 74]. In general, there is a physical world (aka. the physical) and a digital world (aka. the digital). The physical world evolves in time and space. To ensure that the physical system is run in a desired direction, it requires close monitoring, operation, and management. Thus, the role a digital world plays is to model and simulate the dynamics of the physical world in a synchronized fashion, so that the digital can also predict the future states of the physical precisely, which offers a ground for optimal decision making. The physical and the digital exchange data and information flows via a two-way communication. Specifically, the physical sends the data of its own state to the digital, and the digital feeds back the actuation signals to the physical. The actuation would trigger a change in the stage of the physical, and the updated state is sensed and sent back to the digital again. This iterative process runs between the physical and the digital as time unfolds. The sequential states of the physical should move towards a more desired state than without a DT. In other words, the ultimate goal of a DT development is to add values to the physical for improved safety and efficiency. Below, we first formalize transportation DT, and then discuss the relation between DT and cyberphyscial systems.
Definition 2.1
Transportation digital twin (T-DT) is a digital system integrating the pipeline from object detection and tracking, resource allocation, edge-cloud computing and communication, for online simulation, operation, control and management. It is updated online using continuously fed data collected from the physical and send control policies or issue warnings back to the physical, leveraging big data and AI tools. T-DTs are closed-loop with two-way communication, where data, information, and control signals are exchanged with the physical sequentially and reiteratively.
Definition 2.2
Cyber-Physical Systems (CPS) [75] are smart systems that include engineered interacting networks of physical and computational components. CPS holds great potential to enable real-time applications thanks to emerging technologies in sensing, communication, and computing.
A transportation CPS interlinks physical and cyber layers, where the cyber layer consists of sensing, networking, computing, and traffic management application modules (see Fig. 1). The DT encloses the cyber layer, and relies on all the modules for two-way interaction with the physical. To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination. Tab. 1 outlines the comparison of a partial set of related work.
Table 1: Comparison of relevant review papers on DT
Ref Topic Focus Urban Applications CPS role in DT VRU Scope Real Testbeds Integration
Ours Review and position for urban transportation AI-powered CPS architecture Intersection safety warning; Signal control Interlinks physical and cyber layers, DT architecture VRU safety and efficiency A summary of relevant physical testbeds Bidirectional: data exchange, feedback control
[179] Overview for vehicle mobility Connected and automated vehicles (CAV) CAV design and testing (ADAS, ADS) Cloud-based CPS as methodologies Pedestrian safety in CAV testing Mcity, GoMentum Hardware, software, model-in-the-loop
[215] Review for vehicle mobility, architecture, case studies AI-based data-driven cloud-edge-device pipeline Personalized ACC, traffic management, driver modeling DT conceptualized as higher-level evolution of CPS Modeled in physical layer of DT Highway 237 in Mountain View, CA Data and service integration across twins
[89] Review for traffic safety and mobility Hierarchical architecture, cybersecurity Broad transportation safety and mobility services Enables feedback between physical and cyber components As mobile nodes in physical layer, modeled as individual asset for behavior analysis Testbeds mentioned for some literature, but not systematic Integration of data from different levels of twins, and with commun. gateway
[87] Review for smart cities applications Enabling technologies: ML, IoT, CPS, blockchain Transport, water management and automotive technology Facilitates data collection and decision making, DT architecture Ped. safety in Ford’s DT for predictive headlights Durham County Council’s streetlights, IIT’s sewage treatment plant, Insurance Institute for Highway Safety test track Interoperability among different DTs within the smart city ecosystem
[219] Review for transport operation and maintenance Varied levels of adoption across different transport sectors Broad intermodal and ground travel modes CPS enables DT through sensing communication control loop Pedestrian traffic management at train stations N/A Real-time data exchange and synchronization of decision making, integration across platforms
[223] Review for the full lifecycle of transport infrastructure Functions, architectures, integration along the lifecycle Planning and construction, operation and maintenance, and decommissioning and renewal N/A Integrated into an interoperable system for pedestrian pathways, infrastructure, micro-traffic behaviors N/A Data and system integration across various lifecycle stages
The publications on T-DT are normally segmented by different communities and journals, which could prevent researchers from understanding the entire pipeline, from upstream sensing and perception, to downstream transportation applications. For example, [136] on networking DT is primarily focused on the development of communication technology with evaluation on latency, even its use case is transportation. Transportation researchers, who hope to implement this system in real-world, have to seek more details about how traffic dynamics would impact the performance, which is unfortunately missing. This is likely because the authors belong to the society of communication and networking. Another example is that, [118] focused on vehicular DT, heavily rely on the foundational knowledge in CPS, which could be somewhat unfamiliar to many transportation researchers. Realizing such a gap in a large body of literature on T-DT, this paper aims to unify the knowledge by presenting a comprehensive summary upfront, and exemplify the pipeline following the summary with selected applications. After all, the development of a T-DT calls for the interdisciplinary collaboration across transportation, electrical engineering, mechanical engineering, computer science, and human-machine interaction. To foster the readability of the paper, in the next section, we first offer a comprehensive state-of-the-art review on related work along the CPS pipeline.
3 CPS-enabled DT
The development of CPS-enabled DT must engage with expertise in sensing, communication, computing, and human-centric perspectives, where AI methods are backbones.
Figure 1: Schematic diagram of a DT.
3.1 Sensing and perception
Sensors are the “eyes” (and “ears”) of a DT. Tab. 2 summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications. The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
Table 2: Sensors for urban traffic applications (partially adapted from [200, 256])
Sensor Purpose Advantages Disadvantages
Mobile GPS Pedestrian localization 1. Offers global coverage and compatibility with a wide range of devices.
2. Offers reliability and easy access, as it is widely adopted in consumer devices.
1. Accuracy within a few meters, which may be insufficient for safety-critical applications.
2. Poor signals through obstacles.
3. A low update rate for real-time safety applications.
UWB Pedestrian localization 1. Achieves high accuracy, often within a few centimeters, making it suitable for precise indoor and short-range outdoor applications.
2. Operates with low latency, providing real-time updates of position data.
1. Requires anchors to be installed at each corner of the designated area before localization.
2. Operates in a frequency range that may overlap with other wireless technologies, such as Wi-Fi, Bluetooth, or cellular networks [32].
On-board Vehicle Radar Obstacle Detection 1. Outperforms other sensor types at far distances.
2. Detects vehicle speed and position accurately without the need for calibration.
3. Protects privacy, as this sensor type does not record identifiable images of road users.
1. Performs best only when objects move toward or away from the sensor.
2. A Limited number of classes that can be identified due to the lack of color and resolution.
3. Limited field of view when the range is far [170].
Camera Obstacle and lane detection 1. Maintains good resolution when the field of view is wide.
2. Has a long horizon.
1. Difficulties in measuring speed and distance.
2. Performs poorly in bad weather conditions.
Lidar Obstacle detection, 3D mapping 1. Measures distance accurately.
2. Constructs 3D models robustly.
3. Shows promising performance in poor weather.
1. Detects nearby objects poorly [170].
2. Demands high data processing requirements.
3. A shorter effective range than radar.
Accelero -meter Acceleration, driving behavior 1. Detects braking, turning, and accelerating accurately.
2. Integrated with vehicles for behavior analysis [137].
1. Performs poorly for slow or subtle movements.
2. Detects poorly in the presence of noise.
Roadside Infrastructure Camera Object detection at intersections 1. Provides more details, compared to radar and LiDAR, that can be used to differentiate types of vulnerable road users.
2. Covers a larger area where pedestrians are not confined to a narrow path, such as when people crossing midblock [194].
1. Performs poorly in adverse weather conditions.
2. Difficulty in long-term use due to the cost, power supply, and quantity.
3. No privacy protection.
Acoustic Ultrasonic Lane occupancy/vehicle speed 1. Collects data on multiple lanes.
2. Operates during both day and night.
1. Undercounts or overestimates speed.
2. Performs poorly in severe weather.
MMWave Radar Vehicle localization, speed measurement 1. Features a compact size and is easy to install.
2. Offers low latency, within 30ms [212].
3. Penetrates through non-metallic objects.
1. Provides low angular resolution.
2. Measures elevation poorly.
3. Has difficulty with real-time calibration [247, 248].
3.2 Object detection and tracking
Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar. Object tracking is the process of monitoring the detected objects over time to determine their position and movement. Multi-object tracking is concerned with maintaining the identity of the objects and generating their trajectories. Trajectory prediction involves forecasting the future paths of detected and tracked objects. These tasks highly rely on training datasets for urban traffic scenes. Note that there is a much larger size of public datasets collected from on-board vehicles [45], but fewer from other sensor types. Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab. 3.
Table 3: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [45].)
Sensor location Dataset Purpose Sensor Setup Collection region
Infrastructure VIRAT [156], Constellation [197] Urban object detection and tracking, visual event recognition RGB cameras Public outdoor spaces in China, city intersection in New York
Rope3D [242] 2D/3D Road-side object detection, multi-view RGB cameras, LiDAR Streets in Beijing
DeepSense 6G [9] 2D/3D object detection, sensor fusion RGB cameras, mmWave Phase Arrays, LiDAR, Radar Various indoor and outdoor spaces
WILDTRACK [28] Multi-Object Tracking RGB cameras University campus in Zurich
Aerial VisDrone [265], NGSim [155], highD [101], rounD [102], DOTA [228], CitySim [259] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Urban spaces in China and Aachen, intersections in California and Florida
MOT Challenge [40], MOTS20 [203], UAVDT [48] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Various Indoor and Outdoor Scenes
Misc CoCo [121], ADE20K [261], Cityscapes [261] Object detection, semantic segmentation RGB cameras Various indoor and outdoor spaces
Synthetic CARLA [154, 91, 141] Autonomous-driving object detection and segmentation RGB cameras, LiDAR Urban European/North American environments
GTAV [174], Synscapes [143], UrbanSyn [72] Object detection and segmentation RGB cameras, LiDAR Urban European/North American environments
MOTSynth [54] Multi-object tracking RGB cameras Urban European/North American environments
MatrixCity [115] Neural-rendering benchmarking (vehicle/pedestrian-free) RGB cameras Synthetic city environment
Boundless [196] Object detection and segmentation with UE5-synthesised data RGB cameras Synthetic urban environments
Object detection has been studied extensively for urban applications. A large number of studies focus on low-altitude vehicle and pedestrian detection [68, 19]. Many focus on high-altitude aerial environments, where small object detection becomes an important challenge [265, 228, 101, 259]. Single-stage object detectors, following the original single shot multibox detector [130] and You Look Only Once (YOLO) [172] architectures, have become popular due to their real-time deployment capabilities. In the last few years, transformer-based object detection approaches, competitive with YOLO models, have emerged as the state-of-the-art in object detection when designed to be deployed in the real-time setting [22, 266, 267, 258]. Recent progress in YOLO object detection performance has been enabled through multiple small tricks in architecture and training that all together provide significant improvements in empirical performance [205].
When multiple camera views are available, 3D object detection has been studied heavily for autonomous driving [68, 189], as well as for infrastructure-based 3D object detection [156, 197, 242, 9]. Many approaches to 3D object detection use object queries [211], bird’s-eye view transformations [232], or a combination of the two [126].
To build models that make weaker assumptions regarding the sensors, some models have considered the harder task of monocular 3D object detection. MonoCon uses extra regression head branches for learning auxiliary contexts, that are then discarded during inference [132]. DEVIANT is a model architecture equivariant to depth translations [103]. MonoLSS introduces a learnable sample selection module to improve the stability and reliability of the model at test time [117]. Different models have been proposed for infrastructure-based 3D object detection, as many models developed for vehicle-side perception make strong assumptions regarding the position of the cameras. BEVHeight predicts height to the ground to support 3D object detection [234]. CoBEV combines depth and height features to further improve the performance of infrastructure-based 3D object detection [183]. MonoUNI presents the idea of normalized depth, which makes depth prediction independent of camera pitch angle and focal length [94].
To improve the limitations of camera-only perception methods, different sensor combinations are explored. For example, LiDARs or radars, combined with cameras, can detect objects in scenarios where using only the camera is insufficient, such as extreme lighting and weather conditions, or anomalous situations where the camera data is significantly out-of-distribution. Sensor fusion for self-driving cars is now being studied including sensor data for these modalities [29, 123]. These methods often involve the projection of camera, radar and LiDAR features independently to a bird’s-eye view feature space, wherein an aggregation function could be used to merge the features extracted from these different sensors.
Multi-object tracking involves matching newly detected objects with the existing ones by their inter-frame positional and visual similarity information [17, 221]. ByteTrack improves the traditional Hungarian-algorithm-based matching paradigm to gather more comprehensive information [254]. BoT-SORT further incorporates advanced object re-identification modules and a refined Kalman filter for more accurate performances [2]. BoostTrack explores novel distance and shape similarity measurements to deal with ambiguity caused by unreliable detection results [186, 187].
Predicting future trajectories of detected objects is often a crucial part of safety-critical applications. Numerous deep neural network models have emerged as competitive candidates for trajectory prediction over the past few years. Majority of modern architectures for predicting future trajectories of detected objects adopt Recurrent Neural Networks which is responsible for predicting future object positions based on their historical coordinates, together with generative components which handles the variation and flexibility in social interactions [5, 81]. Neural Social Physics model [246] incorporates learnable parameters into explicit physics models built on top of neural networks. SemanticFormer [190] seeks more structural and humanized environment understanding by constructing a semantic knowledge graph. TrajNet++ [100], TDOR [80], and CASPNet++ [schäfer2023caspnet] predict the distributions of future trajectories based on occupancy grid maps. Models like FRM [160] and PPT [122] decompose the prediction task by taking a multi-stage approach. Larger amounts of data comprised of multiple modalities and more comprehensive frameworks have shown increasing importance as is demonstrated by UniTraj++ [57]. Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
3.3 Real-time video analytics
Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges. Video analytics requires the collection and processing of large volumes of video data, which can be resource-intensive and costly. Optimizing computation and network resource usage while maintaining or enhancing the accuracy of analytical results can be challenging. This challenge is further complicated by the need to adapt to varying network conditions, computational resources, and dynamic scene changes in real-time. Tab. 4 provides a comparative analysis of various approaches to address these challenges. The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions. This comparison provided in Tab. 4 highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
Table 4: Comparison of various approaches and optimization objectives in video analytics.
References Approach Optimization objective
SPINN [106], Adaptive offloading [250], Shoggoth [207], Sniper [131], JAVP [236], Auction-base [61] Distributed DNN inference over end devices, edge, and cloud. Optimize throughput, accuracy, and energy consumption under varying network conditions.
CEVAS [233], SAHI [4], CrossRoI [78], Elf [253] Adaptive RoI assignment and frame sampling. Reduce the bandwidth consumption and enhance accuracy.
AdaMask [128], Respire [38], CrossVision [251], VaBUS [206] Leveraging redundant regions on frames and background understanding. Minimize network and computation overhead while ensuring high accuracy.
Elf [253], Mobile edge analytics [120], Sniper [131], JAVP [236], Auction-base [61] Video analytics query scheduling and resource allocation over multiple edge devices. Reduce latency and increase computation resource utilization.
CEVAS [233], Edge-assisted serverless [216] Adaptive model selection. Enhance performance with limited computation resources.
Shoggoth [207], Edge-assisted [99] Online model fine-tuning and model switching. Improve the accuracy and efficiency of real-time video inference on edge devices in changing video scenes.
DAO [149], VaBUS [206], AccMPEG [49], AdaMask [128], ILCAS [224] Adaptive video encoding and compression parameters. Balance low latency, high accuracy, and low compute overhead on edge devices.
AdaDSR [24], AccDecoder [245] Camera-side downsampling and server-side super-resolution upsampling. Balance the trade-offs among accuracy, network cost, and computational cost.
MadEye [222], WiseCam [93] Dynamical orientation adaption of pan-tilt-zoom (PTZ) cameras. Boost the overall accuracy while maintaining the resource cost.
EAIS [239], EALI [238], SERAS [82] Use of an energy-aware scheduler that effectively coordinates batching and dynamic voltage frequency scaling (DVFS) settings. Minimize energy consumption for CNN inference services on high-performance GPUs while meeting latency of Service-Level Objectives.
3.4 Communication and networking
A DT for safety-critical applications requires real-time communications with aggressively low latency. We explore issues related to low latency communications, and survey component technologies and protocols that can be utilized to achieve very low latency.
3.4.1 Real-time requirements and low latency targets
Sensor and control data in a real-time system is subject to latency created by the stages, namely, (1) data acquisition from traffic participants (such as camera recordings and encoding, harvesting data from autonomous vehicles, and collect information from fiber); (2) transmission of data across communications links from sensors to inferencing servers using communications protocols such as Transmission Control Protocol (TCP), Unreliable Data Protocol (UDP) and Real-Time Streaming Protocol (RTSP); (3) data preprocessing (video decoding, and cropping); (4) AI inferencing; (5) higher level reasoning about required feedback to traffic participants; and (6) sending feedback to traffic participants across communications links via low-latency broadcast or dedicated channels over wired and wireless.
Smart city applications can be grouped according to their latency requirements. Many, if not all, pedestrian-associated application (facilitated by pedestrian detection/observations and message notifications) are likely to expect the round trip delay in the range of a couple of seconds. Such latency can be supported by contemporary cameras, communication protocols and inferencing engines. Applications which would attempt to close the observation/notification loop for vehicles moving at about 10 km/h may expect latency in tens of millisecond. Using conventional video compression, RTP/RTSP streaming and edge computing is inadequate to support such latency. This presents the opportunity to pursue novel engineering solutions and research problems.
3.4.2 Communication techniques and protocols
Ultra-Reliable Low Latency Communications (URLLC), a key component of 5G wireless, can help achieve the low latency targets. Along with enhanced mobile broadband (eMBB) and massive machine-type communication (mMTC), URLLC [166] represents one of the three main capabilities of 5G New Radio (5G NR), as standardized by the 3rd Generation Partnership Project (3GPP). In the context of transportation systems, URLLC aims to deliver up to 99.999% reliability and single-digit millisecond latency [198]. However, meeting these performance metrics is challenging in practice due to complex channel environments, particularly in dense urban areas, which can reduce reliability. For intelligent transportation systems, where URLLC may be used as infrastructure backhaul, the target is an end-to-end latency of 30 ms [153].
An emerging technology in this space is Cellular-Vehicle-to-Everything (C-V2X), which has largely replaced the earlier Wi-Fi-based Dedicated Short-Range Communications (DSRC). Unlike DSRC, C-V2X leverages cellular networks, allowing network providers to offer always-on connectivity, which is a critical feature for time-sensitive applications. Additionally, private 5G networks are being developed to ensure this level of connectivity, overcoming the congestion and range limitations inherent to Wi-Fi [53]. In transportation systems, active collaboration between wireless service providers and vehicle manufacturers is in progress to integrate private 5G networks into vehicular networks [201].
Tab. 5 summarizes the key characteristics of commonly used IoT communication protocols, Message Queuing Telemetry Transport (MQTT) [161], Constrained Application Protocol (CoAP) [181], and Hypertext Transfer Protocol (HTTP) [60]. Among them, MQTT emerges as a practical choice, due to its combination of low latency, high scalability, and reliable delivery mechanisms. Its lightweight publish/subscribe model, Quality of Service (QoS) guarantees, and session management make it well suited for real-time data exchange in dynamic, safety-critical environments like urban transportation systems [12, 20].
Table 5: Communication protocol comparison for real-time DT systems.
Protocol Latency Scalability Reliability Best Suited For
MQTT Persistent TCP connection delivers low latency for ongoing messaging Highly scalable with a broker-based publish/subscribe model supporting many-to-many communication across thousands of devices Guarantees reliable delivery with configurable acknowledgment levels, exactly-once delivery via a four-step handshake, persistent sessions Large-scale IoT, real-time sensor/actuator networks
CoAP UDP-based, very low latency when network is stable, but performance degrades with packet loss Limited scalability due to client-server request/response model. Multicast is possible but complex and unreliable at scale Basic two-way acknowledgment, no exactly-once delivery guarantee, no session continuity Resource-limited devices that send small, infrequent data
HTTP Higher latency due to new TCP/TLS handshake and headers per request Limited scalability due to client-server request/response model and verbose ASCII headers Relies on TCP for delivery, with no application-level acknowledgment, retries, or session handling Web APIs, periodic data transfer, backend integration
4 Use cases for urban transportation
The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making. Here we primarily focus on those that are particularly important in urban settings leveraging video analytics, namely, VRU safety warning at road intersections and traffic signal control.
4.1 Safety-time critical applications
Safety-critical applications [6], making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency. Exemplary systems include autonomous driving, Industry 5.0, and mobile and collaborative robots, strictly entailing precise time, computing, communication with bounded latency and AI [23]. Theses systems hold the substantial potential to enhance road safety and save lives. They are, however, risky to test and run, because rare events like automobile collisions are challenging and unethical to replicate. Thanks to emerging technologies in ubiquitous sensing, low-latency high-bandwidth communication, high-speed computing and AI, safety critical applications could potentially be modeled, simulated, processed, and tested in a DT. The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing.
Intersections, where sixty percent of crashes happen [88, 59], are critical bottlenecks of an urban transportation network. To improve urban road safety and increase traffic capacities, safety warning is the key. We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab. 6, and offer an outlook of safety guarantee related methods in Sec. 6.1.3.
Table 6: Literature review on safety warning to VRUs.
Ref Technology Objective Risk Assessment Metrics Evaluation Method/Metric
[157] Sensing: roadside sensors and VRU smartphones
Communication&Networking: 5G, MNO infrastructure, and ITS-G5
Computing: Edge-cloud hybrid computing
Reduces latency and optimizes resource utilization through dynamic service placement. The distance between the VRU and vehicles End-to-end delay: 200 ms
[52] Sensing: Real-time camera detection, YOLOv7
Communication&Networking: 5G, 802.11p, C-V2X, Vehicular Basic Safety Message (BSM)
Computing: Local server equipped with GPU
Develops a video-based vehicular BSMs method with lower error and latency that outperforms the cellular vehicle-to-everything (C-V2X) method. N/A 1. End-to-end delay: <100<100  ms
2. Localization/Speed accuracy
[151] Sensing: GPS on smartphones
Communication&Networking: LTE, Node B, Wi-Fi, Cooperative Awareness Message (CAM)
Computing: CAM server deployed at edge/cloud
Proposes a system using commercial devices and standard messages for road user communication. The distance between VRUs and nearby entities Latency from VRU to CAM <50<50 ms
[139] Sensing: Camera, Android phone’s GPS module
Communication&Networking: 3G, 4G, 5G, and MQTT protocol
Computing: Coral Edge TPU with TensorFlow lite
Develops a traffic safety system using edge computing and 5G to deliver low-latency warnings. The coordinates of pedestrians and cyclists in a driver’s blind spots Latency:
1. 4G - 109.35 ms
2. 5G - 90.95 ms
[119] Sensing: N/A
Communication&Networking: Wi-Fi, C-V2X, 802.11p, ITS-G5, Cooperative Awareness Message
Computing: Edge computing server, smartphones
Develops a system to deliver CAM to VRUs on smartphones using Beacon stuffing without the need for root access to utilize 802.11p. N/A 1. End-to-end latency $\sim$ 2500 ms
2. MAC channel utilization
[191] Sensing: Cameras, Radar, YOLOv3 on NVIDIA JETSON, On Board Units (OBUs)
Communication&Networking: 5G, Fiber, LTE, ITS-G5, and MQTT
Computing: Road side units
Develops a system with sensing and communication, along with fusion and collision detection algorithms, to predict potential collisions and warn VRUs. The distance between VRUs and nearby entities 1. End-to-end latency <300<300 ms
2. Distance error of vehicles and VRUs
[65] Sensing: Camera, real-time object detection
Communication: LTE, Wi-Fi, and MQTT protocol
Computing: GPU on server end
Develops a real-time system with a mobile application to warn pedestrians to avoid vehicle and walker collisions. Time to collision (TTC) 1. End-to-end latency 400 ms.
2. Simulation.
[214] Sensing: CAN, GNSS, roadside, cameras, Google MediaPipe Posekeypoint
Communication&Networking: Fiber, and Simulated V2X
Computing: Server end GPU and portable GPU
Develops a DT framework for connected vehicles and pedestrian in-the-loop simulation. Test it with a V2P collision warning use case. TTC 1. Speed
2. Brake point
3. Distance to conflict point
[69] Sensing: Camera, YOLOv8 object detection.
Communication&Networking: Fiber, LTE, and MQTT protocol
Computing: Server end GPU, resource allocation
Develops a DT pipeline to demonstrate use cases (including intersection safety warning) in urban settings. Distance to the danger zone 1. Localization accuracy
2. Granular latency per frame
4.2 Adaptive traffic signal control
Stochasticity arising from travel demands, un-predicted VRU movement, and traffic gridlocks requires traffic operators to design adaptive traffic signal control (ATSC). With a large amount of agents (including VRUs, traffic lights) continuously interacting in a stochastic environment, classical optimization tools in deterministic environments could fail to capture such complex decision-making processes. There is a surge in employing learning methods, including reinforcement learning and federated learning, to optimize traffic signals [95, 39, 104, 97, 264, 146, 112, 34, 79]. With an increasing concern in data governance and privacy, federated learning witnesses a growing trend that allows optimizing centralized control while preserving distributed data privacy [243, 86, 15, 62, 56, 127]. Related work are summarized in Tab. 7. The last column indicates whether the optimization accounts for the safety of VRUs or not.
Table 7: Summary of papers using (F)RL for ATSC
Ref. RL FL DT Simulator Demand Network VRU
[95] DDPG SUMO
[39] SUMO
[104] A2C Vissim
[97] Vissim
[264] SUMO
[62] A3C FedAvg SUMO
[243] DQN FedAvg Cityflow
[86] PPO FedAvg SUMO
[15] DQN FedAvg SUMO
[146] A2C SUMO
[112] A2C SUMO
[34] A2C SUMO
[79] PPO SUMO
[167] PPO SUMO
[173] D3QN SUMO
[241] DDQN SUMO
Demand: Real-world traffic demand data; Network: Real-world road network.
FedAvg: Federated Averaging, DDPG: Deep Deterministic Policy Gradient, A2C: Advantage Actor Critic, A3C: Asynchronous Advantage Actor Critic, PPO: Proximal Policy Optimization, DQN: Deep Q Learning.
5 Physical Testbeds for DT validation
Physical testbeds are essential for verification, validation, and testing of DTs in diverse and out-of-distribution scenarios. It offers a platform for sensor instrumentation, data collection, actuation, to close the information flow and control feedback loop. Datasets collected from a testbed, if open-sourced, could be used by other researchers who cannot afford building a testbed. Below, we summarize existing testbeds in Tab. 8 based on their types, namely, real-world, closed-track, or scaled-down. Except COSMOS focusing on the dense urban traffic environment, these testbeds are primarily developed for the testing of vehicular technologies and control.
Table 8: Summary of real-world and scaled down testbeds in the U.S.
Type Testbed Open Data Software Hardware Testing scenarios
Real-world COSMOS (NY) [171] software-defined networking (SDN), edge cloud, radio software stack radio access & wireless nodes, mmWave antennas, cameras, GPU/FPGA accelerators V2X networking; traffic sensing; edge computing experiments
Toyota Testbed (CA) [215] AWS cloud platform, traffic simulation platform edge computing, vehicle on-board sensors, full-scale vehicle, onboard computing units, CPU/GPU nodes CAV platooning, ATSC algorithm evaluation, mixed traffic, vehicle control testing
SMTG (NJ) [92] computing cluster, Carla/SUMO co-simulation, Unity visualization lidar, camera, Bluetooth, NVIDIA Jetson AGX Xavier on edge, RSU, Cellular modems real world traffic detection, V2X, DT evaluation, VRU safety.
STARlab (WA) [210] STAR Detection App, VISSIM simulation platform, edge station software V2X Edge Node, Bluetooth and WiFi radios, Raspberry Pi V2X safety, mobility detection
GoMentum (CA) [35] HD Map Database, V2X middleware, sensor middleware full-scale AVs, edge computing, lidar camera, RTK-GPS signalized intersection testing, Construction zone testing, VRU testing
Purdue Testbed (IN) [263] autonomy stack, ROS, cloud-based VLM, whisper, Autoware full-scale AV, lidar, radar, camera, GNSS, high bandwidth networking highway driving, intersection safety, VLM-controlled AV, parking lot testing
Closed track Mcity (MI) [163] unreal virtual Mcity, V2X software stack, driving simulation 32-acre closed test track, perception sensors, RSUs, DSRC, Roadside LiDAR, full-scale AVs ATSC, CAV testing, signalized intersections, AV perception/decision experiments
Scaled down Duckietown (MA) [162] Robot Operating System (ROS) server, signal control framework scaled autonomous vehicle and road, traffic light, Wi-Fi lane following, city navigation, traffic light control
F1Tenth (MA,PA,VA) [14] ROS, Gazebo simulation, RViz visualization, oracle node 1/10 autonomous racecar, NVIDIA Jetson TX2, Wi-Fi, UDP sim-to-real, vehicle racing, vehicle control
Ccity (NY) [168] ROS, self-built simulation, driving emulator 1/8 vehicle with Jetson Orin Nano, camera, UWB localization, traffic signal, Wi-Fi, Raspberry Pi AV lane following, platooning, AV control, traffic signal control, human driving behaviour testing
IDS3C (NY) [26] UDP sockets, IDS 3D city, ROS, Unity engine, Microsoft AirSim plugin 1/25 robotic ground cars, Raspberry Pi, Wi-Fi, UDP, VICON motion capture system CAV control, CPS validation, eco driving, shared mobility, sim-to-real
6 Conclusions and open questions
In this paper, we first review the AI methods applied to every stage of the DT pipeline, from object detection, tracking, prediction, simulation, to traffic operation and management. Then, how DTs are leveraged for use cases including intersection safety warning and traffic signal control are introduced. Following that, the testbeds for DT testing and validation are presented. A growing number of sensors, explosive amounts of data, and increasing computational powers have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline DT pipelines for urban traffic management. Subsequently, we will present emerging trends, challenges, and open questions for the development of DT.
6.1 Emerging trends
While literature on DTs for individual components has been surging, how each element of the DT works collectively and function organically is key to the development of the next-generation DT, which could empower the intelligence and automation of transportation applications.
6.1.1 Engineering the pipeline
Engineering a DT via the integration of multiple subsystems poses technical challenges, and we will name a few.
I. Sensor fusion with time synchronization
Various methods have been proposed to tackle the time-synchronization problem in multi-camera settings for a single intersection or area, either utilizing visual cues or through explicit clock synchronization between the cameras. [188] estimates the spatial transformation between the views. [7] proposes an approach at time synchronization based on the temporal alignment of matching trajectories of entities present in the overlapping scenes. [47] proposes solving a global alignment problem based on video feature descriptors. [16] uses image features to train neural networks for solving the alignment problem. [226] proposes a neural network that uses pose cues to align videos temporally. [184] proposes the use of abrupt lighting changes as temporal cues for facilitating alignment for rolling shutter cameras. Other approaches include estimations of camera capture and transmission latency [107], or clock synchronization [124, 3, 13]. These approaches address the problem of synchronizing sensors across different road intersections. Efficient, low-latency implementation of these methods are vital for synchronization and fusion of camera predictions.
II. Designing edge-cloud architecture and networking
To facilitate the development of a DT in an urban setting with numerous intersections capable of communicating with vehicles and VRUs, an extended network of cloud-connected edge devices and sensors, such as cameras, is required. These sensors generate substantial amounts of real-time data that must be transmitted to edge or cloud systems and processed with bounded latency. The data from all sensors should be integrated into a centralized platform. Given the extensive distribution of these devices, privacy and data security become critical considerations. To safeguard privacy, encoded sensor data is transmitted only to edge devices, where it is processed to extract relevant metadata. Only this metadata is then forwarded to the cloud or central platform for further analysis and integration, ensuring that no raw data or personal information is shared. To this end, data and device federation using federated (reinforcement) learning has gained growing traction [135, 51, 62].
III. Integrated sensing and communication
Integrated sensing and communication (ISAC) is an emerging direction in the design of Beyond-5G wireless networks, enabling transmitted communication waveforms to be opportunistically used as radar-like sensors [125, 220]. In urban mmWave and sub-THz networks, ISAC can enable real-time tracking of vehicles and pedestrians, enhancing sensor fusion algorithms without compromising the network’s primary communication responsibilities. Future research directions may focus on leveraging high angular resolution from densely packed phased arrays to achieve precise beamforming, with the potential to introduce imaging-like capabilities within communication networks [159, 76].
6.1.2 Emerging DT applications
With the increasing demand for urban passenger and goods delivery, emerging technologies like urban sensing, electrification, connectivity and autonomy, and robotics have been gradually transforming urban streetscapes, which pose new challenges to the operation and management of urban infrastructure and public space. DTs are crucial to improve urban safety and mobility in infrastructure planning, service operation and management, and ultimately, policy making.
Here we do not aim to enumerate comprehensive urban transportation applications with DTs, since there exist survey papers [90, 218, 179, 215, 84], which summarize those in operation (e.g., anomaly detection and warning, emergency response), maintenance, and mobility (e.g., transit operation, or driving). Instead, we focus on emerging applications in urban settings, and the potentials the DT holds for them. In Fig. 2, emerging use cases are categorized based on required communication latency (x-axis), spatial resolution (y-axis), and data bandwidth (z-axis), respectively.
Figure 2: Urban T-DT use cases.
6.1.3 AI models
AI methods have been applied to every stage of the DT pipeline. We will point out key challenges of AI methods in transportation.
Physics-informed AI
When it comes to the development of a T-DT, the mutual interaction between domain knowledge and AI is necessary. Domain knowledge in traffic flow and road safety that has been developed for decades, provides valuable insights into every stage of the DT pipeline. In particular, it helps inform technological advancement and testbed deployment, data collection, model selection and training, resource allocation and optimization, as well as performance metrics selection. For instance, the understanding of how traffic evolves across time and space and how entities interact with one another at a road junction, could guide to what degree of granularity and precision the semantic segmentation should be done in object detection and tracking, how communication and networking resources should be prioritized, and what models and metrics should be used for risk prediction.
Safety guarantee of AI
DTs, heavily relying on hardware, software, and algorithms, are vulnerable to risks posed by probabilistic events. Failures in sensors, communication channels, or algorithms can result in erroneous or missing inputs to downstream prediction or decision-making, ultimately compromising the feedback loop from the DT to the physical. The rise of AI-empowered DTs introduces additional risks, including adversarial attacks [262, 41, 225], ethical concerns [71, 192, 204], and trustworthiness in real-world applications [50, 169, 77]. How can DTs be designed with provable safety guarantees, accounting for uncertainties, failures, and attacks? To achieve this, several methods have been proposed, including reachability analysis [10], robust optimization [257], and control barrier/Lyapunov functions [11, 108]. These methods are mathematically rigorous to ensure that DTs generate outputs within safe regimes. With the emergence of AI, adapting these methods to ensure the safety guarantees of AI models is becoming increasingly important. For instance, reachability analysis has been employed in safe reinforcement learning [180, 209]. Robust optimization techniques have also been applied to enhance model robustness towards adversarial attacks [18, 58]. Additionally, control barrier functions and Lyapunov functions have been integrated into neural networks to enforce safety constraints and maintain system stability [66, 27].
Evaluation and validation
As opposed to traditional traffic management systems heavily involved with humans, an AI-powered CPS-enabled traffic management system demands high degree of autonomy, with increased complexity and scale. Since testing traffic management strategies could be unethical and unsafe, DT thus becomes a crucial tool for the test and verification. There does not exist a unified scheme about what to validate, verify, and test in a DT. Here, we would like to decompose this problem into several layers. First, we need to evaluate whether an DT represents its corresponding physical world correctly. Ideally, the digital is expected to be the twin of the physical, accordingly, “Grieves performance test” [73] is a high-level abstract way to compare the difference in the outputs of both the physical and the DT. This is associated with real2sim gap to be defined and elaborated more in the next section.
Second, we need to verify the system behavior of an DT-enabled CPS, and ensure that it performs as desired [185], such as safety, efficiency, accuracy, and timeliness. Such testing is non-trivial, due to the integration of cyber and physical components, as well as their two-way coupling via communication and networking. There are normally four types of tests [185], namely, conformance testing (i.e., whether a system conforms to an expected behavior), robustness testing (i.e., whether a system is robust against stochasticity in environments), fragility testing (i.e., whether the output of a system is robust against perturbation in inputs), and security testing (i.e., whether a system is not affected by cyber-attacks). DTs play important roles in the above tests, while experimental design is crucial to cover comprehensive scenarios leveraging game theory [145] and in recent years, generative models [147].
The presence of human factors along the DT pipeline could complicate the evaluation of cyber-physical-human systems, because of randomness and unpredictability in human behaviors [42]. Humans are not only participants in traffic (as drivers, pedestrians, cyclists), they are also creators and users of DTs. For DTs that involve the feedback to humans, human-in-the-loop test is widely used [244, 113, 42]. Recent years have witnessed a growing trend of using augmented reality and virtual reality to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [214, 105]. Hardware-in-the-loop testing, such as vehicle-in-the-loop [231], could help test the performance of some physical components “live,” while allowing the rest to be simulated within a DT. Since there does not exist a unified approach for the DT testing, evaluation methods, metrics, testing platforms should be developed to facilitate standardized assessment and validation of AI models.
Benchmark datasets & shared testbed resources
Due to limited availability and single perspective of real-world datasets, synthetic data generation from realistic 3D simulators has become increasingly popular for urban computer vision problems. However, open mixed-perspective datasets that adapt to diverse deployment conditions remain scarce. Advances in open-vocabulary object detection offer a promising avenue to address this gap [129, 144, 30].
To advance the application of AI in DT, we must “stand on the shoulders of giants.” In other words, each team does not simply develop one DT for its own internal use. Instead, we hope that an AI-powered DT could be generalized to diverse tasks, transferred to diverse spatiotemporal settings, and shared for co-development among global researchers. Accordingly, we need to standardize application scenarios, benchmark datasets and methods, and unified test environments, for repetitive training and test [43]. Benchmark datasets and methods necessitate performance comparison of any newly proposed AI methods against the state-of-the-art (SOTA) methods. Thus, the transportation community must push to open source data, codes, simulation, algorithms, and results for replicability. For example, standard test platforms, such as Gym, Flower, PettingZoo, and Ray RLlib, are important to benchmark various AI models and methods, which is mostly missing in the transportation community.
While data and methods can often be shared, advancing DTs ultimately depends on shared physical testbeds. These real-world platforms are essential for validation and learning, yet remain costly and space-constrained—calling for new, collaborative approaches from the community. Recent years have seen a growing effort to crowd-source existing testbeds pushed by both agencies and researchers. For example, augmented from the Mcity Test Facility by remote access, Mcity 2.0 [150] enables researchers to connect to the site via a secure cloud interface, run experiments without being physically present, and interact with a mixed-reality environment that overlays virtual actors onto the physical track. In this way, Mcity 2.0 operationalizes standardized, repeatable testing by tightly integrating simulation, physical infrastructure, and remote vehicle control in one unified platform.
6.2 Challenges and open questions
6.2.1 Closing real-to-sim-to-real gap
A key challenge in applying models trained in simulated environments to the real-world is domain shift between the training and test environments. Models trained in simulators might likely experience performance degradation when tested in the real world, where the environment includes unpredictable variations that simulators cannot fully replicate. Such a shift – caused by discrepancies between simulated and real-world conditions – can undermine the generalization of models, resulting in reduced performance in scenarios not represented in training. Accordingly, two gaps exist while establishing a DT, namely, real-to-sim (real2sim) gap (i.e., the deviation of the simulated digital world from the real-world), and sim-to-real (sim2real) gap (i.e., the performance deviation of interventions implemented in the real-world from those simulated in the digital world). The smaller these two gaps are, the closer an DT is to reality.
Although these gaps penetrate through each stage along the DT pipeline, from computer vision, tracking, to prediction, and intervention, sim2real transfer is more studied in object detection [165] and policy learning [164, 178, 134, 36]. The major application area of sim2real transfer in DT is robotic manipulation [1], while that in urban navigation and autonomous driving has witnessed a gradual surge, particularly in computer vision [229] and reinforcement learning [83, 109]. Key methods include domain randomization [193, 147], which introduces variations to augment training data and expose the model to a wide range of scenarios; and domain adaptation [195, 227], a technique of finding mappings to transfer data points observed empirically from two different data distributions.
How do we characterize real2sim gap and control such a gap? This boils down to quantifying errors of the digital representation of a physical world. Minimizing the real2sim gap is key to system identification and representation learning. Depending on observability and internal workings, a system could be modeled as white-box, grey-box, or black-box [98]. Domain knowledge could help represent the physical world with higher accuracy. For example, hybrid twin [33] relies on both data-driven and physics-informed AI. Since calibration of a full model could be time-consuming, expensive, and potentially infeasible, the model reduction philosophy has become popular. Digital shadow [74], a digital model with one-way data exchange from the physical to the digital, takes less effort to build, but could fail to update the state of the physical world once feedback is executed. Digital cousin [37] aims not to build a simulation model that replicates the reality exactly, instead, mainly focus on end-to-end gap, namely, from real-world sensing to intervention.
6.2.2 Prototyping DTs for human intelligence
Many studies define a DT as emerging technologies, namely, object detection and tracking with edge-cloud computing and communication. These technological enablers, however, are essentially “eyes” of a DT, while what really distinguishes a DT from traditional simulators lies in its “brain,” the prediction and decision making modules that are capable of extracting patterns, modeling semantics, and making informed decisions drawing upon what has been seen and perceived. How do we establish a foundational DT that develops human-like intelligence with machine automation? A DT with comparable human-like intelligence or artificial general intelligence (AGI) should consist of a hierarchical cognitive structure analogous to a human’s neural system, backed up by emerging hardware, software, AI algorithms, and API interfaces. The architecture of a proposed DT could consist of:
  1. 1.
    Eyes: object detection and tracking, and perception, powered by convolutional neural networks;
  2. 2.
    Neural systems: edge-cloud networking and computing backbone, powered by resource allocators [111] and cognitive DT [230];
  3. 3.
    Brain: data storage and processing, reasoning and planning, inference and generalization, powered by causal inference and counterfactual analysis [133];
  4. 4.
    Communication and reasoning: natural language processing and vision reasoning, powered by generative AI (GenAI) [182].
A T-DT could be deemed as the world model of a transportation manager. A world model is a mental model learned by an AI agent to simulate the evolution of its environment for action planning and reasoning. It is a special type of DT that relies on the agent’s own sensor information. Moreover, it can be embedded into a DT as the AI agent’s internal, abstract representation of the physical world. A world model emerges from the field of robotic learning, and there is an emerging trend to augment a world model with cognitive and reasoning capabilities for more accurate representation and prediction [255]. Such a trend, we strongly believe, must be the pathway for the next-generation of DTs, despite that DTs could be an external representation of a system that would facilitate engineers to monitor, diagnose, control, and manage the system.
In particular, foundational and GenAI models, which are shown to empower cognitive and reasoning architectures of the world model, have demonstrated great potential in DTs [67, 213, 235]. Large language models (LLM) have been used to generate new data for training [217, 252, 25], enhance interactions between human users and the DT system [249], making personalized recommendations [237, 55] and even automate code generation [46] and creation process of DTs [8]. On the other hand, vision language models (VLM) help augment training datasets for robustness, including generation of critical events [116], videos [64, 63] and simulations [202], as well as visual question answering [208, 70]. Thus, we believe that GenAI-powered DT will be the next-generation of DTs for efficient and safe traffic management.
6.2.3 Limits in AI
Despite the promising future of AI-powered DTs, application of AI to DTs could face challenges. As opposed to classical statistical methods, AI algorithms are generally black boxes where their inner workings are not transparent to developers nor users. Thus, interpretability or explainability using Shapley value [140], PIDL [44], symbolic regression [142], or Kolmogorov-Arnold networks [138] can potentially reveal to some degree the rationale underlying the predictions. Unlike humans, AI models do not understand causes and heavily rely on correlations to make predictions. Without knowing causality could render AI methods incapable of generalizing to unseen data. Thus, augmenting AIs with causal reasoning and inference could increase its deductive capabilities [175, 176, 177]. In addition, generative AIs could produce hallucination, which might lead to nonphysical predictions or unrealistic decision making. Introducing physics based domain knowledge could help fine tune these models, enhance inductive biases, and generate more meaningful outputs. On the other hand, the emergence of LLM could facilitate the alignment of AI models with human preference. For example, reinforcement learning from human feedback [96] has seen a rapid growth for preferential learning in autonomous driving [85], as well as the generation of more realistic traffic DTs [21].
Last but not the least, there has been a growing trend in developing safe AI systems aligned with human values and objectives. For example, recent studies have focused on evaluating LLMs in terms of toxicity [31], privacy [240], ethics [158], and fairness [148], indicating that LLMs are not sufficiently safe. It is thus crucial to continually achieve safety guarantees of AI, especially for the implementation of DTs.
In a nutshell, transportation applications in urban settings are generally challenging to design, develop, deploy, and test for their potential unsafe and unethical consequences. Thus, AI-empowered DT plays a critical role in effective implementation of these applications. Despite relatively sparse literature in this domain, we review an ensemble body of literature on how to leverage emerging technologies in sensing, communication, edge and cloud computing, for urban traffic management. We hope this paper can serve as a pointer to help researchers and practitioners understand SOTA methods and gaps on the development of DTs; a bridge to initiate conversations across interdisciplinary researchers; and a road map to exploiting potentials of DTs for urban transportation applications.
Acknowledgments
This work was supported by NSF CPS-2038984 and ERC-2133516. We thank Mengxuan Liu for her assistance in generating the 3D models.